Field and CPL Updates

2008-2009 Antarctica

January 25, 2009
Update provided by Bruce H. Vaughn
project update image
Dr. Ken Taylor, WAIS Divide Chief Scientist, Dr. Julie Palais, NSF-OPP Glaciology Program Manager, and Mr. Bruce Vaughn, SCO Field Rep, outside the arch facility at WAIS Divide. Photo: Dr. Julie Palais

The 2008-2009 field season is coming to a close and we have reached our goals. Here is a summary of recent activities.

Drilling operations:
Late in the evening of January 22nd we ceased drilling for the 2008-2009 field season at depth of 1511.9 meters. The stopping point was predetermined by the maximum amount of ice that we could safely store in the basement of the arch on carts for the winter. A total of 931.9 meters of core was processed, measured for DEP, and is now in the basement. The preliminary DEP data indicate that the last ice drilled this season is approximately 7693 ybp in age. John Fegyveresi cut and prepared 9 samples for physical properties taken every ~20 meters beginning at 1320 meters depth, and are also stored in the basement. The new long cable for the drill remains McMurdo, despite numerous attempts at scheduled flights (see list below) to get it here in time to be spooled by ICDS on the winch. The cable that was used this year has been spooled off the winch, and when the new cable arrives (likely after most of the ICDS and SCO crew have departed) it will be staged inside the drilling arch for re-spooling onto the winch at the beginning of next season.

Other Activities:
All of camp has been packing up for the last 3 days. Both the drill arch and core-processing arch have been prepared for the winter. The AC units have been winterized, and work areas have been cleaned, organized and inventoried for next year. ISC boxes have been staged inside the arch, and more to come from MCM will also go in the arch. Later, incoming carpenters will complete the job by adding necessary plywood and tarps to protect the arch ends from blowing winter snows. Science Jamesway was taken down today. Winter berms are built and being loaded with cargo.

With any luck we have two primary flights scheduled for WAIS Divide tomorrow January 27th, and we'll send out science cargo (vessel bound), and nearly all SCO and ICDS personnel along with visitors that arrived during this last week. A small group of camp staff will remain to completely break down winterize the camp. Much cargo is palletized and is waiting to be retrograded to McMurdo.

Fuel on site: 5,500 gallons. More is expected before close of season.

Received: (1) LC 130 (D029 1/19)
Canceled: (5) due to weather as WAIS Divide: (Sk 23, Sk 52, Sk 61) Cancelled: (2) due to no airplanes: (Sk 43, Sk 53)
Boomerang: (1) Sk 34 -- circled for two hours before returning to MCM

Passenger movements: Monday Jan. 19th flight.
IN: NSF: Dick Armstrong (turn-around),T-350: Alexander Shturmakov; BFC: Leah Biezun and Rebekah Foy; Planners: Matthew Kippenhan, Chad Naughton, Rob Edwards; EHS: Michael Stone; Sci Cargo: Brian Connell.
OUT: RPSC - Victoria (Nurse); I-151 Sylvain Masclin; I-205 -(4) - Srihdar Anandakrishnan, Leo Peters, Huw Horgan, Joe Macgregor; I-188 -(1)- Anthony Hoch.

It's the end of a successful field season, and while many chores remain to be done both here and in McMurdo, I think many people are already looking to life after WAIS Divide in New Zealand and beyond.